Please ;') A right triangle is a triangle with one angle measuring 90°. In a right triangle, the sides are related by Pythagorean Theorem, where c is the hypotenuse (the side opposite the 90° angle). Find the hypotenuse when the other 2 sides’ measurements are 3 feet and 4 feet.
Answer by sarah_adam(201)

First lets state the Pythogeran Theorem:
If a,b and c are the sides of a right angled triangle where c is the hypotenuse(i.e., the side opposite to the 90° angle ) then according to the Pythogeran Theorem
a^2+b^2= c^2
here in this problem a = 3 ft and b = 4 ft
so substituting the values of a,b and c in the equation we get
3^2+4^2 = c^2
9+16 = c^2
25 = c^2
therfore c = sqrt(25)
which implies c = 5 ft